# Enhancing LLMs for Power System Simulations: A Feedback-driven Multi-agent Framework

### v2

@@ I Introduction
The challenges outlined above can be grouped into three main limitations: (i) limited simulation-specific knowledge, (ii) restricted reasoning capabilities for simulation tasks, and (iii) imprecision in function and option application. Enhancing the simulation capability of LLMs requires addressing these limitations. However, few existing works have explicitly focused on overcoming the above barriers to improve the simulation capability of LLMs, even though LLM applications in power systems are growing rapidly. Specifically, in power systems, LLMs have been recently used to translate language-based rules into mathematical constraints to facilitate optimal power flow (OPF) analysis, bridging the gap between rule-based and computational methods [8]. Researchers have also leveraged LLMs to interpret decision-making processes in real-time market, enhancing the transparency of deep reinforcement learning systems by revealing decision rationales [9]. Additionally, LLMs have been employed to retrieve and summarize documents in response to specific power system queries [10]. In other work, LLMs have been applied to derive OPF solutions iteratively by utilizing historical cost-solution data [10]. LLMs have also facilitated gathering user preference w.r.t. electric vehicles charging, where user inputs are integrated to refine functions for EV charging optimization problems [10]. Furthermore, by combining LLMs with retrieval-augmented generation (RAG), researchers develop a carbon footprint accounting system capable of dynamically retrieving and integrating real-time, domain-specific carbon data [11]. Moreover, for cybersecurity applications, LLMs have played a role in anomaly detection to enhance system security [12]. In forecasting, LLMs, such as LLaMa2, have been used to integrate social event data into time series models, enhancing the accuracy and contextual relevance of predictions, for example, in electricity demand [13]. Also, LLMs have supported perception analysis by evaluating media sentiment and public acceptance levels of solar power initiatives, providing insights into public opinion trends [14]. Moreover, a comprehensive benchmarking framework for LLMs has been proposed for the energy domain in [15], helping to establish standardized evaluation criteria for LLMs in energy-related applications. On the other hand, potential cybersecurity threats, arising due to the application of LLMs in power systems, have also been analyzed and summarized [16].
@@
  • It proposes an enhanced RAG module with an adaptive query planning strategy and a triple-based structure (i.e. linking options, functions, and their dependencies) for the knowledge base. This module expands the LLM’s accessible knowledge in an efficient and cost-effective manner. Also, it enables LLMs to better identify and interpret simulation functions, options, and their logical relationships than the standard RAG.
@@ II-A Adaptive Query Planning
Remark: The adaptive query planning strategy goes well beyond simple keyword extraction. It employs semantic recognition via few-shot CoT, which can translate descriptions into relevant keywords. This enables the retrieval agent to infer implicit simulation functions and options from context—even when explicit terms are missing. For example, consider the request: “For the IEEE 24-bus Reliability Test System, solve the AC optimal power flow (OPF) problem with the de-commitment of expensive generators,” which is the initial request in complex task 7 for MATPOWER (see Section V for more details). Rather than extracting generic phrases like “AC optimal power flow” or “de-commitment,” the retrieval agent—guided by few-shot CoT prompting for function mapping—directly identifies “runuopf” as the keyword, i.e., the specific MATPOWER function for OPF with de-commitment.
@@ II-B Triple-based Structure Design for Knowledge Base
To overcome these issues, this paper proposes an additional, easy-to-construct retrieval repository: a triple-based structured option document. The following details the approach to build such a document for a given simulation tool:
@@
Note that the inclusion of the function dependency in the document enables retrieval for logical relationships. As will demonstrated in case studies, this supplementary repository significantly enhances retrieval efficiency and improves the accuracy of simulation code generation by preserving the logical context.
@@ III Enhanced Reasoning Module
The enhanced RAG module provides LLMs with retrieval results tailored to a simulation request. However, it is still essential to strengthen the LLM’s reasoning abilities. This ensures the generation of correct simulation codes based on the retrieval results. Hence, a coding agent is required (i.e., another LLM) to write codes for simulations tasks. This agent needs to fully understand its role, assigned tasks, reasoning path, and contextual knowledge, including retrieval results, when handling simulation tasks.
@@ III-B Reasoning Framework
Each of these actions is further clarified with tool-specific coding examples in the prompt. While the examples are tool-dependent, the rest of the framework remains general. Overall, the above reasoning framework highlights again that the key to handling simulation tasks: correctly identifying and combining functions and options.
@@ III-C Knowledge Integration
  • Dynamic Retrieval Knowledge: Complements static knowledge by incorporating real-time, request-specific details. This is achieved through the integration of the aforementioned RAG module, which retrieves relevant information on option formats, values, and function dependencies, for accurate code generation.
@@
Eventually, by integrating both static and dynamic knowledge, as well as the above structured reasoning framework, the coding agent is expected to generate accurate code to address the simulation request.
@@ III-D Adaptability
The enhanced reasoning module leverages few-shot chain-of-thought prompting techniques, which embody universal principles and remain tool-agnostic. Additionally, supported by the RAG module, the reasoning module integrates two levels of knowledge: the static and dynamic knowledge. The static basic knowledge provides fundamental, coarse-grained insights—such as the basic principles of using a simulation tool. Meanwhile, dynamic retrieval knowledge supplies detailed, fine-grained information. This dual-layer approach is universally applicable across different tools. As a result, the enhanced reasoning module adapts seamlessly to a variety of simulation tasks and environments.
@@ IV Environmental Acting Module with Feedback
Despite the reinforcement brought by the enhanced RAG and reasoning modules, the coding agent may still encounter errors during simulation code generation. To address this, it is essential to enable direct interaction between the LLM and the simulation environment. It allows the agent to receive execution feedback and iteratively refine its code. To this end, this paper proposes an environmental acting module with an error-feedback mechanism that integrates with both the RAG and reasoning modules, as illustrated in Fig. 5. The components of this module are described in the following.
@@ IV-B Error Handling and Feedback Loop
TABLE I: Evaluated schemes by distinct combinations of the proposed strategies within the framework
@@ IV-D Adaptability
The environmental acting module features a tool-independent design. First, it leverages the error detection and reporting systems common to many power system simulation tools. Second, it formulates the error-feedback loop in a general manner, as detailed in Section IV-B. Each component is independent of any specific simulation tool. This design not only supports automatic error correction but also improve adaptability to diverse simulation platforms.
@@ IV-E Pseudocode for the Full Framework
Remark: It is important to emphasize that the proposed framework is built on the premise that simulation tools are learnable by humans through sufficient documentation. If a human can use a simulation tool based on its manuals, then the proposed integrated RAG, reasoning, and feedback modules can replicate and enhance this process to improve efficiency. Hence, for scenarios where documentation is absent, it is suggested to generate necessary documents, possibly with LLMs’ help. These documents can then be incorporated into the proposed framework.
V Case Study
The following sections detail the case study configurations, followed by an analysis of the simulation outcomes for Daline and MATPOWER. Eventually, the cost of using LLMs to perform power system simulations is discussed. All supporting materials, including prompts, knowledge bases, simulation tasks, training datasets, and results (i.e., the generated codes and their benchmarks, totaling 870 coding files) will be available upon the acceptance of this paper.
@@ V-A Settings
TABLE II: Configuration settings for evaluations (for the LLM model versions, please see Table I or the captain of each evaluation figure/table)
@@
  • $P_{t,i}=100$ points if the simulation result is exactly correct with no irrelevant settings in the code.
  • $P_{t,i}=50$ points if the simulation result is correct but contains irrelevant settings.
  • $P_{t,i}=0$ points if the simulation result is incorrect.
@@
In the following experiments, $N_{\max}^{t}=3$ ($\forall t$) is set for Daline tasks, and $N_{\max}^{t}=5$ ($\forall t$) is set for MATPOWER tasks, reflecting the increased complexity of the latter.
@@ V-B Evaluation on Daline
The high success rate of 93.13% over all taks for GPT4o-Full is due to the cumulative effects of using the complete proposed framework. Comparing other schemes with GPT4o-Full gives an indication of the impact of omitted strategies. For instance, although GPT4o-NP includes most reasoning enhancement strategies, it lacks the triple-based structured option document, resulting in a reduced success rate of 81.37%. On the other hand, omitting the few-shot CoT for reasoning, as in GPT4o-NCS, lowers success to 65.19%. When few-shot CoT is employed, but the proposed query planning is omitted, as in GPT4o-RSR, the success rate for complex tasks drops to 66.67%, particularly due to the deteriorated performance for the complex tasks. Similar comparisons can be drawn across all schemes. However, the goal is not to determine which strategy provides the highest improvement. Instead, this paper aims to emphasize that high success relies on the combined effect of multiple strategies.
@@
In general, more complex tasks — those with multiple sub-requests — tend to increase the likelihood of errors in LLMs, resulting in lower success rates compared to standard tasks, as observed across most schemes. However, with the proposed full framework, GPT4o-Full, the performance gap between complex and standard tasks narrows significantly, as shown in both Fig. 7 and Fig. 8. This suggests that, with enhanced reasoning capabilities and the more effective RAG design, GPT4o-Full effectively identifies and addresses the sub-requests within complex tasks, similar to how it handles standard tasks. This enables LLMs to better manage complex tasks.
Figure 7: Distribution of scores achieved across attempts for each evaluated scheme (simulation environment: Daline; GPT4o version: gpt-4o-2024-05-13).
Figure 8: Success rates for each scheme, itemized by all tasks combined, complex tasks only, standard tasks only, as well as for the first attempt success rate and the final attempt success rate (simulation environment: Daline; GPT4o version: gpt-4o-2024-05-13).
@@
Enhancing error-reporting systems in simulation tools is beyond this paper’s scope. However, for tools with ambiguous error messages, the following is proposed:
@@ V-C Evaluation on MATPOWER
The evaluation results on MATPOWER are illustrated in Figs. 9, 10, and 11. Specifically, Fig. 9 presents the scores achieved by each evaluated scheme in individual attempts when managing complex tasks. Fig. 10 shows the distribution of scores across attempts for each scheme, and Fig. 11 depicts the success rates of each scheme. The outcomes observed here align closely with the results on Daline. This alignment allows us to focus primarily on comparative analyses across schemes in the subsequent discussion.
@@
Despite MATPOWER being a widely-used and well-documented tool with extensive resources available online, the latest high-performance LLMs, such as GPT4o and o1-preview (renowned for its reasoning capability), struggle to perform simulations reliably. For instance, both GPT4o-Sole and o1p-Sole show a 0% success rate on complex tasks, and GPT4o-Sole achieves only 27.77% success on standard tasks. Even with RAG and the whole knowledge base, GPT4o-SR reaches a success rate of only 13.75% for complex tasks and 52.96% for standard tasks. In contrast, the fully equipped framework, GPT4o-Full, achieves a remarkable 96.85% success rate across all tasks, with a breakdown of 93.75% on complex tasks and 97.77% on standard tasks, underscoring the framework’s effectiveness.
@@ V-D Comparison with Supervised Fine-tuning
The underperformance of GPT4o-Sole-SFT could stem from the representativeness of the training datasets—despite significant human effort in generating them, these datasets may not and cannot capture every nuanced detail of MATPOWER. To rigorously eliminate the possibility that the limited dataset is responsible for the performance gap, an additional evaluation was conducted using the Daline simulation environment. In this experiment, the full testing dataset (i.e., the Daline simulation tasks used for evaluation) was used as the SFT training dataset. Specifically, all 34 testing simulation tasks, augmented with 16 additional tasks, were employed for fine-tuning, yielding a training dataset of 32,346 tokens. GPT4o-Sole-SFT was fine-tuned for 3 epochs, with a batch size of 1 and a learning rate multiplier of 2 (again, all automatically configured). The score distributions and specific success rates are presented in Fig. 13 and Table III. This evaluation shows that GPT4o-Sole-SFT improves the success rate of GPT4o-Sole from 0% to 51.96% overall, with 28.57% for complex tasks and 58.03% for standard tasks. Yet, even with the entire testing dataset used for fine tuning, GPT4o-Sole-SFT still fails to reach high accuracy.
@@
V-E Cost Analysis

### v3

@@ I Introduction
The challenges outlined above can be grouped into three main limitations: (i) limited simulation-specific knowledge, (ii) restricted reasoning capabilities for simulation tasks, and (iii) imprecision in function and option application. Enhancing the simulation capability of LLMs requires addressing these limitations. However, few existing works have explicitly focused on overcoming the above barriers to improve the simulation capability of LLMs, even though LLM applications in power systems are growing rapidly. Specifically, in power systems, LLMs have been recently used to translate language-based rules into mathematical constraints to facilitate optimal power flow (OPF) analysis, bridging the gap between rule-based and computational methods [8]. Researchers have also leveraged LLMs to interpret decision-making processes in real-time markets, enhancing the transparency of deep reinforcement learning systems by revealing decision rationales [9]. Additionally, LLMs have been employed to retrieve and summarize documents in response to specific power system queries [10]. In other work, LLMs have been applied to derive OPF solutions iteratively by utilizing historical cost-solution data [10]. LLMs have also facilitated gathering user preference w.r.t. electric vehicle charging, where user inputs are integrated to refine functions for EV charging optimization problems [10]. Furthermore, by combining LLMs with retrieval-augmented generation (RAG), researchers develop a carbon footprint accounting system capable of dynamically retrieving and integrating real-time, domain-specific carbon data [11]. Moreover, for cybersecurity applications, LLMs have played a role in anomaly detection to enhance system security [12]. In forecasting, LLMs, such as LLaMa2, have been used to integrate social event data into time series models, enhancing the accuracy and contextual relevance of predictions, for example, in electricity demand [13]. Also, LLMs have supported perception analysis by evaluating media sentiment and public acceptance levels of solar power initiatives, providing insights into public opinion trends [14]. Moreover, a comprehensive benchmarking framework for LLMs has been proposed for the energy domain in [15], helping to establish standardized evaluation criteria for LLMs in energy-related applications. On the other hand, potential cybersecurity threats, arising due to the application of LLMs in power systems, have also been analyzed and summarized [16].
@@
  • It proposes an enhanced RAG module with an adaptive query planning strategy and a triple-based structure (i.e., linking options, functions, and their dependencies) for the knowledge base. This module expands the LLM’s accessible knowledge in an efficient and cost-effective manner. Also, it enables LLMs to better identify and interpret simulation functions, options, and their logical relationships than the standard RAG.
@@ II-A Adaptive Query Planning
Remark 1: The adaptive query planning strategy goes well beyond simple keyword extraction. It employs semantic recognition via few-shot CoT, which can translate descriptions into relevant keywords. This enables the retrieval agent to infer implicit simulation functions and options from context, even when explicit terms are missing. For example, consider the request: “For the IEEE 24-bus Reliability Test System, solve the AC optimal power flow (OPF) problem with the de-commitment of expensive generators,” which is the initial request in complex task 7 for MATPOWER (see Section V for more details). Rather than extracting generic phrases like “AC optimal power flow” or “de-commitment,” the retrieval agent—guided by few-shot CoT prompting for function mapping—directly identifies “runuopf” as the keyword, i.e., the specific MATPOWER function for OPF with de-commitment.
@@ II-B Triple-based Structure Design for Knowledge Base
To overcome these issues, this paper proposes an additional, easy-to-construct retrieval repository: a triple-based structured option document. The following details the approach of building such a document for a given simulation tool:
@@
Note that the inclusion of the function dependency in the document enables retrieval for logical relationships. As will be demonstrated in case studies, this supplementary repository significantly enhances retrieval efficiency and improves the accuracy of simulation code generation by preserving the logical context.
@@ III Enhanced Reasoning Module
The enhanced RAG module provides LLMs with retrieval results tailored to a simulation request. However, it is still essential to strengthen the LLM’s reasoning abilities. This ensures the generation of correct simulation codes based on the retrieval results. Hence, a coding agent is required (i.e., another LLM) to write code for simulation tasks. This agent needs to fully understand its role, assigned tasks, reasoning path, and contextual knowledge, including retrieval results, when handling simulation tasks.
@@ III-B Reasoning Framework
Each of these actions is further clarified with tool-specific coding examples in the prompt. While the examples are tool-dependent, the rest of the framework remains general. Overall, the above reasoning framework highlights again that the key to handling simulation tasks is: correctly identifying and combining functions and options.
@@ III-C Knowledge Integration
  • Dynamic Retrieval Knowledge: Complements static knowledge by incorporating real-time, request-specific details. This is achieved through the integration of the aforementioned RAG module, which retrieves relevant information on option formats, values, and function dependencies for accurate code generation.
@@
Eventually, by integrating both static and dynamic knowledge, as well as the above-structured reasoning framework, the coding agent is expected to generate accurate code to address the simulation request.
@@ III-D Adaptability
The enhanced reasoning module leverages few-shot chain-of-thought prompting techniques, which embody universal principles and remain tool-agnostic. Additionally, supported by the RAG module, the reasoning module integrates two levels of knowledge: static and dynamic knowledge. The static basic knowledge provides fundamental, coarse-grained insights, such as the basic principles of using a simulation tool. Meanwhile, the dynamically-retrieved knowledge supplies detailed, fine-grained information. This dual-layer approach is universally applicable across different tools. As a result, the enhanced reasoning module adapts seamlessly to a variety of simulation tasks and environments.
@@ IV Environmental Acting Module with Feedback
Despite the reinforcement brought by the enhanced RAG and reasoning modules, the coding agent may still encounter errors during simulation code generation. To address this, it is essential to enable direct interaction between the LLM and the simulation environment. It allows the agent to receive execution feedback and iteratively refine its code. To this end, this paper proposes an environmental acting module with an error-feedback mechanism that integrates with both the RAG and reasoning modules, as illustrated in Fig. 5. The components of this module are described below.
@@ IV-B Error Handling and Feedback Loop
TABLE I: Evaluated schemes by distinct combinations of the proposed strategies; a checkmark indicates inclusion of the corresponding strategy.
@@ IV-D Adaptability
The environmental acting module features a tool-independent design. First, it leverages the error detection and reporting systems common to many power system simulation tools. Second, it formulates the error-feedback loop in a general manner, as detailed in Section IV-B. Each component is independent of any specific simulation tool. This design not only supports automatic error correction but also improves adaptability to diverse simulation platforms.
@@ IV-E Pseudocode for the Full Framework
Remark 2: It is important to emphasize that the proposed framework is built on the premise that simulation tools are learnable by humans through sufficient documentation. If a human can use a simulation tool based on its manuals, then the proposed integrated RAG, reasoning, and feedback modules can replicate and enhance this process to improve efficiency. Hence, for scenarios where documentation is absent, it is suggested to generate necessary documents, possibly with LLMs’ help. These documents can then be incorporated into the proposed framework.
V Case Study
The following sections detail the case study configurations, followed by an analysis of the simulation outcomes for Daline and MATPOWER. Eventually, the cost of using LLMs to perform power system simulations is discussed. All supporting materials, including prompts, knowledge bases, simulation tasks, training datasets, and results (i.e., the generated codes and their benchmarks, totaling 870 coding files) will be openly available upon acceptance.
@@ V-A Settings
TABLE II: Configuration settings for evaluations (for the LLM model versions, please see Table I or the caption of each evaluation figure/table)
@@
  • $P_{t,i}=100$ points if all requirements are fully satisfied and no irrelevant settings appear in the code.
  • $P_{t,i}=50$ points if all requirements are satisfied, but the code includes irrelevant or unnecessary settings that do not affect task completion.
  • $P_{t,i}=0$ points if any requirement is not satisfied.
@@
In the following experiments, $N_{\max}^{t}=3$ ($\forall t$) is set for Daline tasks, and $N_{\max}^{t}=5$ ($\forall t$) is set for MATPOWER tasks, reflecting the increased complexity of the latter. In addition, the scoring was conducted manually by human experts during the evaluation process.
Remark 3: To further clarify what constitutes the aforementioned irrelevant settings, preliminary examples are provided here for illustration, with full details to be shown in the subsequent case studies. These include: (i) explicitly setting parameters to their default values when not required (e.g., in DALINE, setting data.baseType as TimeSeriesRand in Standard Task 2, and data.program as acpf in Standard Task 7, both by GPT4o-PR); and (ii) including unnecessary function calls, such as using define_constants in MATPOWER when no constants are actually used (e.g., Standard Task 27 by GPT4o-Sole). Such settings reflect imperfections in code generation and therefore receive partial credit.
@@ V-B Evaluation on Daline
The high success rate of 93.13% over all tasks for GPT4o-Full is due to the cumulative effects of using the complete proposed framework. Comparing other schemes with GPT4o-Full gives an indication of the impact of omitted strategies. For instance, although GPT4o-NP includes most reasoning enhancement strategies, it lacks the triple-based structured option document, resulting in a reduced success rate of 81.37%. On the other hand, omitting the few-shot CoT for reasoning, as in GPT4o-NCS, lowers success to 65.19%. When few-shot CoT is employed, but the proposed query planning is omitted, as in GPT4o-RSR, the success rate for complex tasks drops to 66.67%, particularly due to the deteriorated performance for the complex tasks. Similar comparisons can be drawn across all schemes. However, the goal is not to determine which strategy provides the highest improvement. Instead, this paper aims to emphasize that high success relies on the combined effect of multiple strategies.
@@
In general, more complex tasks (i.e., those with multiple sub-requests) tend to increase the likelihood of errors in LLMs, resulting in lower success rates compared to standard tasks, as observed across most schemes. However, with the proposed full framework, GPT4o-Full, the performance gap between complex and standard tasks narrows significantly, as shown in both Fig. 7 and Fig. 8. This suggests that, with enhanced reasoning capabilities and the more effective RAG design, GPT4o-Full effectively identifies and addresses the sub-requests within complex tasks, similar to how it handles standard tasks. This enables LLMs to better manage complex tasks.
Figure 7: Distribution of scores achieved across attempts for each evaluated scheme. For the definitions of scheme notations (e.g., GPT4o-Full), refer to Table I; the notations apply hereafter. (simulation environment: Daline; GPT4o version: gpt-4o-2024-05-13).
Figure 8: Success rates for each scheme. “All” refers to the aggregated success rate across all tasks; “All-Complex” and “All-Standard” report the success rates calculated exclusively for complex and standard tasks, respectively. “First Attempt” represents the success rate achieved on the first attempt, before any error correction, and “Final Attempt” reflects the success rate after all permitted attempts are completed. The same interpretation applies hereafter (simulation environment: Daline; GPT4o version: gpt-4o-2024-05-13).
Figure 9: Distribution of feedback loops (attempts) required for successful task completion in complex (left) and standard (right) tasks. (simulation environment: Daline; GPT4o version: gpt-4o-2024-05-13).
@@
V-B5 Distribution of Attempts for Task Completion
To further examine the feedback mechanism, Fig. 9 presents the distribution of attempts (feedback loops) required to successfully solve complex and standard tasks. For complex tasks, baseline schemes (e.g., GPT4o-Sole) frequently fail or require multiple attempts, indicating that resolving complicated queries (i.e., tasks with more sub-queries and dependencies) poses significant challenges. In contrast, GPT4o-Full performs markedly better, with most tasks completed within one or two attempts. For standard tasks, all schemes achieve improved results, and failures are infrequent. GPT4o-Full again shows clear advantages, solving the majority of standard tasks in a single attempt. Even less-equipped schemes perform relatively well on standard tasks, suggesting that these tasks impose lower demands on reasoning, retrieval, and error correction, and are thus more manageable for LLMs.
Remark 4: Enhancing error-reporting systems in simulation tools is beyond this paper’s scope. However, for tools with ambiguous error messages, the following is proposed:
@@ V-C Evaluation on MATPOWER
The evaluation results on MATPOWER are illustrated in Figs. 10, 11, and 12. Specifically, Fig. 10 presents the scores achieved by each evaluated scheme in individual attempts when managing complex tasks. Fig. 11 shows the distribution of scores across attempts for each scheme, and Fig. 12 depicts the success rates of each scheme. The outcomes observed here align closely with the results on Daline. This alignment allows us to focus primarily on comparative analyses across schemes in the subsequent discussion.
@@
Despite MATPOWER being a widely-used and well-documented tool with extensive resources available online, the latest high-performance LLMs, such as GPT4o and o1-preview (renowned for the reasoning capability), struggle to perform simulations reliably. For instance, both GPT4o-Sole and o1p-Sole show a 0% success rate on complex tasks, and GPT4o-Sole achieves only 27.77% success on standard tasks. Even with RAG and the whole knowledge base, GPT4o-SR reaches a success rate of only 13.75% for complex tasks and 52.96% for standard tasks. In contrast, the fully equipped framework, GPT4o-Full, achieves a remarkable 96.85% success rate across all tasks, with a breakdown of 93.75% on complex tasks and 97.77% on standard tasks.
@@
Figure 13: Distribution of feedback loops (attempts) required for successful task completion in complex (upper) and standard (lower) tasks. (simulation environment: MATPOWER; GPT4o version: gpt-4o-2024-05-13; o1p-Sole is only tested by complex tasks).
V-C3 Distribution of Attempts for Task Completion
Compared to Fig. 9, a similar pattern of the attempt distribution is observed under the MATPOWER simulation environment, as shown in Fig. 13. GPT4o-Full still maintains a clear advantage by completing the majority of tasks in the first or second attempt. These results further confirm the effectiveness and generalizability of the proposed framework across different simulation environments.
@@ V-D Comparison with Supervised Fine-tuning
Figure 16: Scores achieved by each evaluated scheme in individual attempts when handling tasks with different difficulty levels; the bar on the right represents the score scale, where 100, 50, and 0 indicate different score values (simulation environment: MATPOWER; GPT4o version: gpt-4o-2024-08-06).
The underperformance of GPT4o-Sole-SFT could stem from the representativeness of the training datasets—despite significant human effort in generating them, these datasets may not and cannot capture every nuanced detail of MATPOWER. To rigorously eliminate the possibility that the limited dataset is responsible for the performance gap, an additional evaluation was conducted using the Daline simulation environment. In this experiment, the full testing dataset (i.e., the Daline simulation tasks used for evaluation) was used as the SFT training dataset. Specifically, all 34 testing simulation tasks, augmented with 16 additional tasks, were employed for fine-tuning, yielding a training dataset of 32,346 tokens. GPT4o-Sole-SFT was fine-tuned for 3 epochs, with a batch size of 1 and a learning rate multiplier of 2 (again, all automatically configured). The score distributions and specific success rates are presented in Fig. 15 and Table III. This evaluation shows that GPT4o-Sole-SFT improves the success rate of GPT4o-Sole from 0% to 51.96% overall, with 28.57% for complex tasks and 58.03% for standard tasks. Yet, even with the entire testing dataset used for fine-tuning, GPT4o-Sole-SFT still fails to reach high accuracy.
V-E Evaluation under Increased Attempt Budget
To further analyze how tasks evolve toward either successful completion or termination by the stopping criterion, Fig. 16 presents the scores achieved in individual attempts with an extended attempt budget of up to 50. Meanwhile, task difficulty is explicitly characterized by the number of sub-queries within each task. This analysis offers a more comprehensive view of how different schemes perform under varying task complexities, especially when granted extensive opportunities for error correction.
The results in Fig. 16 reveal a clear distinction between schemes and task difficulty levels. Baseline methods struggle significantly with complex tasks — they often either reach the maximum iteration limit (i.e., the stopping criterion) or terminate without execution errors but still fail to satisfy all sub-queries. Typical failure reasons include: (i) invalid option names, which prevent proper execution and remain unresolved even after 50 feedback loops (e.g., GPT4o-Sole-SFT in the 2nd and 5th tasks, GPT4o-Sole in the 5th task, and GPT4o-SR in the 5th task); (ii) correct option names with invalid parameter values produce incorrect but executable code (e.g., GPT4o-Sole in the 2nd to 4th tasks and GPT4o-SR in the 3rd task); (iii) missing critical option settings, which result in early termination without errors but produce incorrect outputs (e.g., GPT4o-PR in the 5th task).
These observations confirm that repeated error-feedback loops alone are insufficient for convergence on challenging tasks without strong reasoning and retrieval capabilities. In contrast, GPT4o-Full exhibits robust performance across all difficulty levels, completing all tasks — including those with numerous sub-queries — within only a few attempts. Notably, GPT4o-Full only failed at the first attempt for the 5th task, where it misused an option name. This error, however, was successfully corrected in the second attempt.
V-F Cost Analysis